ROTATIONAL SPECTRA AND HYPERFINE STRUCTURE FOR A TITANIUM SANDWICH COMPLEX, C_{5}H_{5}TiC_{7}H_{7}^{a}

\underline{STEPHEN KUKOLICH}, CHAKREE TANJAROON and KRISTEN KECK, Department of Chemistry, University of Arizona, Tucson, AZ 85721.

Microwave spectroscopy measurements and density functional theory calculations are reported for the cyclopentadienylcycloheptatrienyltitanium complex, C_{5}H_{5}TiC_{7}H_{7}. This appears to be the first microwave work on a complex containing the cycloheptatrienyl ligand. Rotational transition frequencies for this symmetric-top complex were measured in the 4-13 GHz range using a Flygare-Balle-type pulsed beam spectrometer. The spectroscopic constants obtained for the normal isotopomer are B= 771.78907(38), D_{J}= 0.0000295(41), and D_{JK}= 0.001584(73) MHz. The quadrupole hyperfine splittings for C_{5}H_{5}^{48}TiC_{7}H_{7} were clearly observed, and eQq_{aa}= 18.432(90) MHz. Analysis of the rotational constants indicates that bond lengths in the gas phase are about 0.02Å longer than those reported for the solid-state X-ray structure. The calculated Ti-C bond lengths are shorter for the C_{7}H_{7} ligand (r(Ti-C)=2.21Å) than for the C_{5}H_{5} ligand (r(Ti-C)=2.34Å), and the C_{7}H_{7} H atoms are displaced 0.15Å out of the C7 plane, toward the Ti atom.

aSupported by THE NATIONAL SCIENCE FOUNDATION CHE0304969